An XYZ business can currently produce 251 per hour of 5 amp fuses. The supplier claims that a new machine has been purchased and installed that will boost the manufacturing rate. The mean hourly production on the new machine was 259, with a sample standard deviation of 6 per hour, according to a sample of ten hours randomly selected from last month.


1
Expert's answer
2021-10-08T08:09:08-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq251"

"H_1:\mu>251"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

"df=n-1=10-1=9" degrees of freedom, and the critical value for a right-tailed test is "t_c =1.833113."

The rejection region for this right-tailed test is"R = \{t: t > 1.833113\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{259-251}{6/\sqrt{10}}\approx4.21637"

Since it is observed that "t = 4.21637 > 1.833113=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for right-tailed, "\alpha=0.05, df=9, t=4.21637" is "p =0.001126," and since "p=0.001126<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 251, at the "\alpha = 0.05" significance level.

Therefore, there is enough evidence to claim that the new machine is faster, at the "\alpha = 0.05" significance level.